# The Feasibility of Topic-Based Watermarking on Academic Peer Reviews

### v1

@@ Abstract
Large language models (LLMs) are increasingly integrated into academic workflows, with many conferences and journals permitting their use for tasks such as language refinement and literature summarization. However, their use in peer review remains prohibited due to concerns around confidentiality breaches, hallucinated content, and inconsistent evaluations. As LLM-generated text becomes more indistinguishable from human writing, there is a growing need for reliable attribution mechanisms to preserve the integrity of the review process. In this work, we evaluate topic-based watermarking (TBW), a lightweight, semantic-aware technique designed to embed detectable signals into LLM-generated text. We conduct a comprehensive assessment across multiple LLM configurations, including base, few-shot, and fine-tuned variants, using authentic peer review data from academic conferences. Our results show that TBW maintains review quality relative to non-watermarked outputs, while demonstrating strong robustness to paraphrasing-based evasion. These findings highlight the viability of TBW as a minimally intrusive and practical solution for enforcing LLM usage in peer review.
@@
Alexander Nemecek, Yuzhou Jiang, Erman Ayday Case Western Reserve University {ajn98, yxj466, exa208}@case.edu
@@ 1 Introduction
As large language models (LLMs) continue to evolve, their adoption has accelerated, particularly in academic writing Dergaa et al. (2023); Editorials (2023). LLMs are widely used for language polishing, literature search, and low-novelty writing, often producing text nearly indistinguishable from human-authored content. Many conferences now explicitly permit authors to use LLMs for low-novelty tasks, provided that authors retain full responsibility for the content ACL (2025a); NeurIPS (2025); ICML (2025a). These policies uphold pre-LLM expectations around authorship and accountability while adapting to new technological norms.
@@
In contrast, the use of LLMs by peer reviewers is widely prohibited ACL (2025b); NeurIPS (2025); ICML (2025b). Such practices risk confidentiality breaches, low-quality evaluations, and data exposure to third-party systems Zhou et al. (2024); Maini et al. (2024). Recent empirical studies suggest, however, that LLM-assisted reviews are already present in major conferences, leading to inflated scores, reduced reviewer confidence, and distortions in paper rankings Liang et al. (2024); Latona et al. (2024); Ye et al. (2024). These findings underscore the urgency of developing attribution mechanisms to detect and manage unauthorized LLM usage.
@@
As LLM-generated content increasingly mirrors human writing, distinguishing between machine- and human-authored reviews has become difficult. Stylistic cues alone are insufficient for reliable attribution, especially in the absence of disclosure Mitchell et al. (2023). This creates an urgent need for technical mechanisms to trace the provenance of peer reviews. A widely explored approach is watermarking, which has been adopted across various domains to embed imperceptible, machine-detectable signatures into generated text Zhao et al. (2024). Recent methods, such as topic-based watermarking, bias generations toward semantically aligned tokens that are robust and minimally intrusive. However, existing work focuses on general-domain text, with limited analysis in peer review Liu et al. (2024); Zhao et al. (2023).
In this paper, we present the first focused evaluation of topic-based watermarking in the context of academic peer reviews. Rather than proposing a new algorithm, we apply an existing lightweight, topic-guided watermarking scheme to this domain-specific, policy-sensitive task. Topic-based watermarking (TBW) offers a balance of efficiency, robustness to paraphrasing, and minimal impact on generation quality, making it suitable for peer review, where stylistic fidelity and semantic coherence are critical. It also supports domain adaptation through customizable topic lists, aligning well with the structured topical nature of peer reviews. Moreover, TBW relies on a topic-matching assumption that naturally holds in this setting, where reviews are expected to stay aligned with the subject of the paper.
Our goal is to assess whether TBW can preserve review quality and semantic fidelity while offering reliable attribution under realistic adversarial settings. We evaluate across three LLM configurations: a pretrained base model, a few-shot configuration, and a fine-tuned model using authentic reviews from AI and ML conferences. Our analysis examines generation quality, semantic preservation, classifier-based attribution, and robustness to paraphrasing. We further compare TBW against general-purpose watermarking methods and find that TBW offers better preservation of text quality, highlighting its suitability for domain-sensitive tasks like peer review.
@@
Without effective attribution mechanisms, the credibility and rigor of academic conferences could erode, leading to lower-quality evaluations and increased reliance on potentially unverifiable, machine-generated feedback. Watermarking provides a practical and minimally disruptive approach for LLM accountability, helping to safeguard academic standards while accommodating the evolving role of generative models.
2 Related Work
Since the release of ChatGPT, LLMs have been rapidly adopted across various stages of the academic workflow. Their use has raised concerns about authorship and peer review integrity. Most conferences and journals now permit authors to leverage LLMs; however, this permissive stance does not extend to peer reviewers. Leading venues such as NeurIPS and ACL explicitly prohibit the use of LLMs by reviewers NeurIPS (2025); ACL (2025b). These policies reflect growing concerns around review quality, including the risk of shallow or hallucinated feedback, reduced technical depth, and breaches of confidentiality that would compromise the double-blind review process Li et al. (2024).
Despite these restrictions, recent studies suggest that LLM-assisted reviews are already present at major conferences. Liang et al. (2024) estimate that 5–15% of reviews were substantially modified using LLMs, with affected reviewers showing lower confidence and less engagement during rebuttals. Latona et al. (2024) report similar trends and observe a score inflation effect, while Ye et al. (2024) show that even subtle LLM manipulations can shift paper rankings. Together, these findings underscore the risks unauthorized LLM use poses to peer review fairness and rigor.
Given the increasing use of LLMs for peer review generation, recent work has focused on detecting and attributing such content. Much of this research explores classifier-based detection or semantic similarity methods aimed at identifying AI-generated text. For example, Yu et al. (2025) propose a detection method based on the semantic similarity between a known LLM-generated review and a test review, flagging a review as machine-generated when similarity exceeds a threshold. Similarly, Kumar et al. (2025) introduce a partition-based method under the assumption that a review contains both human- and LLM-written components. They segment the review into distinct points, complete each segment with a reference LLM, and measure semantic similarity between these completions and the original text to detect potential LLM involvement.
However, these detection methods fail under paraphrasing or hybrid-review scenarios, where even minor edits or partial human rewriting can evade detection. To address this limitation, watermarking offers a promising alternative by embedding identifiable signals directly into the generated text. One foundational method is the KGW algorithm Kirchenbauer et al. (2023), which partitions the model’s vocabulary into “green” and “red” token sets. During generation, the model is subtly biased to sample more frequently from the “green” list, which acts as a watermark-carrying set, while avoiding tokens in the “red” list. This results in output text that biases outputs toward “green” tokens with minimal quality loss. Variants aim to improve robustness and preserve quality Liu et al. (2024); Zhao et al. (2023); Hou et al. (2024).
More recently, commercial systems have also entered this space. For example, Google’s SynthID-Text watermarking system employs a strategy called Tournament Sampling, in which candidate tokens are ranked according to randomized watermarking functions, and the highest-ranked token is selected during generation Dathathri et al. (2024). While both academic and commercial watermarking approaches have shown promise, they are primarily evaluated on general-purpose domains such as news or encyclopedic text, and rarely tested under the stylistic and ethical constraints found in peer review.
While a few frameworks target peer review watermarking Rao et al. (2025), they rely on tightly integrated pipelines and lack evaluation across adaptation modes. Topic-based watermarking (TBW) Nemecek et al. (2024), originally proposed for open-domain text, provides a lightweight, semantically guided alternative. We adapt TBW to peer review by aligning token selection with domain-relevant topics, preserving generation quality while supporting practical reviewer attribution. Section 3.2 details this adaptation.
3 Methodology
Our goal is to evaluate the applicability of topic-based watermarking in the domain of academic peer review. We investigate whether such watermarking can preserve the quality and semantic integrity of generated reviews, while enabling robust attribution under paraphrasing attacks. We describe our data collection, model configurations, watermarking integration, and evaluation procedures.
3.1 Peer Review Generation Task
We simulate realistic LLM-based peer review generation by training and prompting language models to write reviews conditioned on a paper’s title and abstract. We use the abstract rather than the full paper because full submissions often exceed typical context window limits and are less readily available in structured form. This section describes the dataset used, the model variants we examine, and our prompting and fine-tuning strategies.
3.1.1 Dataset
To evaluate topic-based watermarking in the context of peer review, we compile a dataset of paper titles, abstracts, and corresponding reviews from ICLR and NeurIPS conferences using the OpenReview API OpenReview (2024). Each review includes a summary, strengths and weaknesses, and a final recommendation score. To minimize the risk of including LLM-generated reviews, we restrict our dataset to conferences held before the public release of ChatGPT (November 2022) OpenAI (2022). Specifically, we collect reviews from ICLR 2018–2023 and NeurIPS 2021–2022, noting that the ICLR 2023 review phase, despite the conference date, occurred prior to ChatGPT’s availability ICLR (2023). Although language models existed before this, they were not widely adopted in peer review workflows at scale.
The final dataset contains approximately 19,000 reviews. For each paper, we randomly sample a single review to construct prompt-completion training pairs, ensuring diversity in reviewer perspectives while avoiding overrepresentation of any one submission. Detailed review counts by conference are provided in Appendix A.1.
3.1.2 Model Configurations
To assess the feasibility of topic-based watermarking across varying levels of model adaptation and reviewer effort, we utilize the Llama-3.1-8B Grattafiori et al. (2024) open-source language model in three configurations: base, few-shot, and fine-tuned. The base configuration uses the pretrained model without any additional training or prompt engineering, simulating minimal reviewer effort. The few-shot setting provides the model with example peer reviews as part of the input prompt, enabling it to better replicate the expected format and tone with lightweight guidance. Finally, the fine-tuned configuration involves additional supervised training on peer review data using parameter-efficient methods, resulting in a model that is more aligned with the review-writing task and capable of generating coherent, domain-adapted outputs. This model size offers a practical balance between computational efficiency and generation quality, making it suitable for experiments involving multiple training configurations.
3.1.3 Prompting and Few-shot Learning
Title: [TITLE] Abstract: [ABSTRACT] Please write a detailed review.
3.1.4 Fine-tuning Setup
For fine-tuning, we follow a supervised instruction-tuning setup where each instance consists of an input prompt (title + abstract + instruction) and a target completion (review text). The dataset is split into training (80%), validation (10%), and test (10%) subsets. We fine-tune using LoRA (Low-Rank Adaptation) with 4-bit quantization, enabling gradient checkpointing and early stopping. The objective is to improve the fluency and consistency of generated reviews while approximating the tone and structure typical of human-written peer reviews. Fine-tuning hyperparameters, model setup, and training procedure details are provided in Appendix A.2.
3.2 Topic-Based Watermarking
@@
Topic-based watermarking (TBW) Nemecek et al. (2024) is a semantic-aware watermarking method that subtly influences a language model’s token selection process to leave a detectable signature. Unlike earlier schemes such as KGW Kirchenbauer et al. (2023), which rely on randomly partitioned vocabularies, TBW constructs topic-specific token subsets (“green lists”) aligned with the semantic content of the input prompt. This design helps preserve fluency and coherence while enhancing robustness against paraphrasing and token-level edits. We briefly summarize the TBW generation and detection process as applied in our setup.
3.2.1 Token-to-Topic Mappings
TBW first assigns tokens to topic-specific green lists using semantic similarity. A small set of generalized topics ${t_{1},\dots,t_{K}}$ is defined, each represented by an embedding $\mathbf{e}_{t_{i}}$ computed via a sentence embedding model. Each token $v\in V$ in the model’s vocabulary is embedded as $\mathbf{e}_{v}$, and its cosine similarity with each topic embedding is calculated:
$\text{sim}(v,t_{i})\;=\;\frac{\mathbf{e}_{v}\;\cdot\;\mathbf{e}_{t_{i}}}{\|\mathbf{e}_{v}\|\;\|\mathbf{e}_{t_{i}}\|}.$
If the maximum similarity exceeds a threshold $\tau$, the token is assigned to the green list $G_{t_{i}}$ for the most similar topic. Tokens that do not meet this threshold are placed in a residual set and evenly distributed across all green lists to maintain full vocabulary coverage.
While the original implementation used general-purpose topic categories (e.g., technology, sports), we adapt the topic set to align with the thematic structure of academic reviews, better reflecting the linguistic and topical distribution of this domain.
3.2.2 Generation
Once topic-specific green lists are defined, TBW applies a watermark by biasing the model’s output distribution during generation. For each input prompt, the most relevant topic is identified using a lightweight keyword extraction method (e.g., KeyBERT). If the extracted topic exactly matches one of the predefined topic labels, the corresponding “green” list is selected. If no exact match is found, topic embeddings are computed and the most similar predefined topic is selected based on cosine similarity.
At each decoding step, the model produces a probability distribution over its vocabulary $V$. TBW modifies this distribution by adding a small logit bias $\delta$ to all tokens in the selected green list. This increases the likelihood of sampling topic-aligned tokens after applying the softmax function, subtly guiding the generation process without altering the model architecture or requiring multiple decoding passes. The watermark strength is controlled by the value of $\delta$: higher values produce stronger attribution signals but cause detectable shifts in word choice or token distribution. The approach is model-agnostic and incurs minimal overhead, making it compatible with standard generation pipelines.
3.2.3 Detection
TBW uses a statistical test to detect whether a given text contains a watermark. Detection mirrors the generation process by recovering the relevant topic inferred from the input text using the same keyword or embedding-based matching procedure, and the corresponding green list $G_{t^{*}}$ is recovered.
@@ 3.2.3 Detection
The number of green-list tokens $g$ is then counted in the text $\mathbf{z}_{\text{test}}$, and compared to the total number of tokens $n$. A $z$-score quantifies whether the green-token rate exceeds an expected baseline proportion $\gamma$:
@@
If $z>z_{\text{threshold}}$, the text is classified as watermarked. The threshold can be tuned to balance sensitivity and specificity, and the method is prompt- and model-agnostic at inference time, requiring access only to the generated output. Importantly, the detection process is model-agnostic and does not require access to the model logits or original input prompt.
3.2.4 Watermarking Configurations
To ensure consistency with the original TBW implementation while adapting it to the domain of peer review, we retain most of the original parameter settings. We use the same sentence embedding model, all-MiniLM-L6-v2 Reimers and Gurevych (2020), to encode tokens and topic labels into a shared semantic space. Topic extraction from input prompts is performed using KeyBERT Grootendorst (2020), as in the original work.
Following the TBW framework, we partition the vocabulary into green lists based on semantic similarity to a predefined set of $K=4$ topics. While the original implementation used general-purpose topics such as {animals, technology, sports, medicine}, we adapt these categories to reflect the structure and content of machine learning conference reviews. Specifically, we define the following domain-specific topics: {theory, applications, models, optimization}. These topics are designed to capture broad themes in peer review content from venues like ICLR and NeurIPS, and can be adjusted to suit different research domains.
We apply a logit bias of $\delta=2.0$ to green-list tokens during generation, consistent with values reported in prior literature Kirchenbauer et al. (2023). For token-to-topic assignment, we primarily use a cosine similarity threshold of $\tau=0.7$, but also evaluate a lower threshold of $\tau=0.3$ to assess how watermark detection and text quality vary under relaxed alignment constraints.
3.3 Rationale for Topic-Based Watermarking
While several general-purpose watermarking methods exist, we select topic-based watermarking (TBW) for its unique combination of robustness, adaptability, and minimal performance overhead. Prior work has demonstrated that TBW is resilient to paraphrasing, while preserving generation quality and incurring no additional inference cost Nemecek et al. (2024). This property is important in the peer review setting, where paraphrasing represents a realistic threat model where a reviewer seeking to obscure LLM use may rewrite or rephrase parts of a generated review, but is unlikely to introduce noise or degrade the review’s usefulness or semantic integrity. These full-paraphrase attacks, rather than token-level perturbations or synthetic distortions, reflect plausible reviewer behavior under current policy constraints.
TBW’s semantic token-level biasing strategy is well-suited to this context. It subtly steers generation toward topic-consistent vocabulary without disrupting fluency or style, both of which are critical in high-stakes peer review writing. In addition, TBW supports domain adaptation through customizable topic lists, and relies on a topic-matching assumption that naturally holds in peer review, where content is expected to stay aligned with the paper under evaluation.
Finally, the peer review task inherently satisfies TBW’s core assumption of topic consistency between the prompt and the generated output. One known limitation of TBW is the Topic Matching Assumption, which requires that the generated text remain semantically aligned with the prompt topic. In general-purpose settings, this assumption can be violated due to topic drift or open-ended generation. In peer review, however, this risk is minimal, as the input (e.g., paper title and abstract) directly constrains the review content. A reviewer cannot reasonably produce a review on a different topic than the paper itself. As such, TBW aligns naturally with the structural and semantic constraints of the peer review task.

### v2

@@ Abstract
Large language models (LLMs) are increasingly integrated into academic workflows, with many conferences and journals permitting their use for tasks such as language refinement and literature summarization. However, their use in peer review remains prohibited due to concerns around confidentiality breaches, hallucinated content, and inconsistent evaluations. As LLM-generated text becomes more indistinguishable from human writing, there is a growing need for reliable attribution mechanisms to preserve the integrity of the review process. In this work, we evaluate topic-based watermarking (TBW), a semantic-aware technique designed to embed detectable signals into LLM-generated text. We conduct a systematic assessment across multiple LLM configurations, including base, few-shot, and fine-tuned variants, using authentic peer review data from academic conferences. Our results show that TBW maintains review quality relative to non-watermarked outputs, while demonstrating robust detection performance under paraphrasing. These findings highlight the viability of TBW as a minimally intrusive and practical solution for LLM attribution in peer review settings. [^1] Our code and data is available at https://github.com/ANCP2021/Watermarking-LLM-Conference
@@
Alexander Nemecek1*,   Yuzhou Jiang1,   Erman Ayday1 1Case Western Reserve University {ajn98*, yxj466, exa208}@case.edu
@@ 1 Introduction
As large language models (LLMs) continue to evolve, their adoption has accelerated in academic writing dergaa2023human; editorials2023tools. LLMs are widely used for language polishing, literature search, and low-novelty writing. Many conferences now explicitly allow authors to use LLMs for certain tasks, provided that authors retain full responsibility for the content emnlp2025cfp; neurips2025llm; icml2025cfp. These policies uphold pre-LLM expectations around authorship and accountability while adapting to new technology.
@@
In contrast, the use of LLMs by peer reviewers is widely prohibited emnlp2025review; neurips2025llm; icml2025review. Such practices risk confidentiality breaches, low-quality evaluations, and data exposure zhou-etal-2024-llm; maini2024llm. Recent empirical studies suggest, however, that LLM-assisted reviews are already present in major conferences, leading to inflated scores, reduced reviewer confidence, and distortions in paper rankings liang2024monitoring; latona2024ai; ye2024we. These developments underscore the urgency of developing attribution mechanisms to detect and manage LLM usage in peer review.
@@
Distinguishing between machine- and human-authored reviews has become difficult, as LLM-generated content continues to improve. This creates an urgent need for technical mechanisms to trace review provenance. Watermarking offers a promising approach, embedding imperceptible, machine-detectable signatures into generated text zhao2024sok. However, existing work focuses on general-domain text, with limited analysis in peer review contexts liu2024a; zhao2023provable.
In this paper, we present the first systematic evaluation of topic-based watermarking (TBW) in the context of academic peer reviews, a high-stakes domain with distinct structural, semantic, and ethical constraints that differentiate it from general-purpose text generation settings. While TBW was originally proposed for open-domain text nemecek2024topic, its applicability to specialized domains with strict quality requirements and adversarial threat models has remained unexplored. We address this gap by adapting TBW to peer review workflows and conducting a comprehensive assessment of its feasibility as a practical attribution mechanism for conference organizers.
Our contribution is threefold. First, we demonstrate that peer review represents an ideal use case for TBW due to its inherent structural properties: reviews must remain topically aligned with the paper under evaluation, satisfying TBW’s topic-matching assumption naturally, unlike open-ended generation tasks where topic drift undermines detection. Second, we introduce domain-specific adaptations including custom topic sets ({theory, applications, models, optimization}) aligned with machine learning conference reviews, and we systematically evaluate TBW across three model configurations: base, few-shot, and fine-tuned, representing varying levels of model adaptation. Third, we assess robustness under realistic paraphrasing threats (PEGASUS and DIPPER), reflecting scenarios where reviewers may rephrase generated content to improve clarity or evade detection while preserving review quality.
@@
Without effective attribution mechanisms, the credibility and rigor of academic conferences could erode, leading to lower-quality evaluations and increased reliance on potentially unverifiable, machine-generated feedback. Watermarking provides a practical and minimally disruptive approach for LLM accountability, helping to safeguard academic standards while accommodating the evolving role of generative models. Related work on LLM watermarking and detection approaches is provided in §A.
2 Topic-Based Watermarks
@@
Topic-based watermarking (TBW) nemecek2024topic is a semantic-aware watermark that subtly influences a language model’s token selection process to leave a detectable signature. Unlike earlier schemes such as KGW kirchenbauer2023watermark, which rely on randomly partitioned vocabularies using a secret seed, TBW constructs topic-specific token subsets (“green lists”) aligned with the semantic content of the input prompt through a deterministic, keyless mapping. This approach helps preserve fluency while enhancing robustness against paraphrasing.
Token-to-Topic Mappings. TBW assigns tokens to topic-specific green lists using semantic similarity without requiring a master key or secret seed. A small set of topics $t_{1},\dots,t_{K}$} is defined, each represented by an embedding $\mathbf{e}_{t_{i}}$ computed via a sentence embedding model. Each token $v\in V$ in the model’s vocabulary is embedded as $\mathbf{e}_{v}$, and its cosine similarity with each topic embedding is computed. If the maximum similarity exceeds a threshold $\tau$, the token is assigned to the green list $G_{t_{i}}$ for the most similar topic. Tokens that do not meet this threshold are placed in a residual set and evenly distributed across all green lists to maintain full vocabulary coverage. This deterministic assignment process is reproducible given only the embedding model, topic set, and threshold $\tau$ as no secret parameters are involved. [^2] While this keyless design is suitable for practical threats like paraphrasing, TBW can be extended with a keyed residual assignment (e.g., using a secret seed to distribute non-matching tokens) if cryptographically stronger adversaries are of concern.
Generation & Detection. During generation, the most relevant topic is identified from the input prompt using keyword extraction, and TBW adds a small logit bias $\delta$ to all tokens in the corresponding green list. This increases the likelihood of sampling topic-aligned tokens while preserving the model architecture and generation efficiency. The watermark strength is controlled by $\delta$: higher values produce stronger attribution signals but may cause detectable shifts in word choice.
@@
For detection, TBW recovers the relevant topic from the input text and counts green-list tokens $g$ relative to total tokens $n$. A $z$-score quantifies whether the green-token rate exceeds an expected baseline proportion $\gamma$:
@@
If $z>z_{\text{threshold}}$, the text is classified as watermarked.
Rationale for TBW. We select TBW for its combination of robustness, adaptability, and minimal overhead nemecek2024topic. TBW is particularly well-suited to peer review contexts where paraphrasing represents a realistic threat model where reviewers may rephrase generated content to improve clarity or avoid detection, but are unlikely to introduce noise that would degrade quality.
Additionally, TBW’s requirement of topic alignment between the input prompt and generated text naturally holds in peer review, where content must remain semantically aligned with the paper under evaluation. This structural constraint provides an advantage over general-domain watermarking because the detector (conference organizer) has direct access to the same paper submission used at generation time, they can reliably recover the correct topic and reconstruct the identical green lists without requiring any secret key. This eliminates key dependency entirely, the deterministic token-to-topic mapping combined with guaranteed topic alignment means detection requires only the public submission text, not hidden parameters.
Finally, TBW’s semantic biasing strategy preserves fluency and style while supporting domain adaptation through customizable topic sets $t_{1},\dots,t_{K}$} that can be tailored to specific research fields or venues.
3 Experimental Setup
We simulate realistic LLM-based peer review generation by training and prompting models to write reviews conditioned on a paper’s title and abstract. We condition on paper abstracts rather than full text due to context length constraints and structured data availability.
3.1 Dataset
We construct a dataset of paper titles, abstracts, and corresponding reviews from ICLR and NeurIPS conferences using the OpenReview API openreview-api. To minimize risk of including LLM-generated content, we restrict data to conferences before ChatGPT’s release (November 2022) openai_chatgpt_2022: ICLR 2018-2023 and NeurIPS 2021-2022. The final dataset contains approximately 19,000 reviews, each including summary, strengths/weaknesses, and recommendation scores. For each paper, we randomly sample a single review to construct prompt-completion pairs, ensuring reviewer diversity while avoiding overrepresentation. Detailed statistics are in §C.1.
3.2 TBW Domain Adaptation
We adapt TBW to peer review by modifying the topic sets while retaining original parameter settings. Instead of general-purpose topics (e.g., technology, sports), we define domain-specific topics: {theory, applications, models, optimization} to capture themes in ML conference reviews. We use logit bias $\delta=2.0$ and similarity threshold $\tau=0.7$ for primary experiments, with additional evaluation at $\tau=0.3$. Complete parameter details are provided in §C.2.
3.3 Model Configurations
To assess the feasibility of TBW across varying levels of model adaptation and reviewer effort, we utilize the pretrained Llama-3.1-8B base checkpoint grattafiori2024llama in three configurations: base, few-shot, and fine-tuned. The base configuration uses the pretrained model checkpoint directly without additional training or prompt engineering, simulating minimal reviewer effort. The few-shot setting provides the same base model with example peer reviews as part of the input prompt, enabling it to replicate the expected format and tone with lightweight guidance. The fine-tuned configuration involves supervised training on peer review data using parameter-efficient methods, resulting in a model more aligned with the review-writing task. This model size offers a practical balance between computational efficiency and generation quality suitable for our multi-configuration experiments. Detailed setup parameters for few-shot prompting and fine-tuning are provided in §C.3.